OpenStudy (anonymous):

x-y=2 y=4-2x solve system using substitution please show me steps i did this like 7 times and didnt get it right

OpenStudy (radar):

You will substitute for either the x or the y. In this case the y would be the best choice as it is already in the form for substitution.

OpenStudy (radar):

Taking the first equation we substitute for the y and will now have; x-(4-2x)=2. Now we remove the parenthesis being careful with signs. x-4+2x = 2 Combine similar terms getting 3x-4 =2 Now add 4 to both sides of the equal sign resulting in 3x=6 Dividing both sides by 3 solves for x giving: x=2 Now we solve for y by plugging in the value 2 for x, I choose the 2nd equation. y=4-2(2) y=4-4 y=0

OpenStudy (anonymous):

ohhhh thankkk youuuu i didnt put parenthesis so i got confused

OpenStudy (radar):

You are welcome, those signs preceding parentheses can be a bear!